OpenStudy (anonymous):

HELPP!!! log(10^6 x^2)= A. 12 B. 6+2log(x) C. 12log(x) D. 6x^2

zepdrix (zepdrix):

\[\log(ab)=\log(a) + \log(b)\]\[\log(c^d)=dlog(c)\] Hmm ok we need to use a couple rules of logarithms to figure this one out :D So for the first part, let a=10^6 b=x^2 What would your problem look like after you apply the first rule i listed? :D

OpenStudy (anonymous):

log(10^6)(x^2)=log(10^6)+log(x^2)

zepdrix (zepdrix):

k cool c: Now we want to apply that other rule, we'll look at each log separately. Let c=10 d=6 And for the other one, c=x d=2